OpenStudy (anonymous):

would anyone know what the definition of transforming a quadratic equation is, and a real life example

OpenStudy (accessdenied):

I imagine that transforming a quadratic is the same as for any function. It would be a shift or compression/stretch horizontally or vertically of the quadratic equation. You might show y = x^2 transformed into y = A(x - h)^2 + k Shifted h units right, k units up, stretch/compression factor A. As for an example, how about the path of a ball kicked in the air? What happens if you kicked it on a platform 5 ft above the ground? Or kicked the ball to the left or right of the original position? Maybe kicked it with more or less impact as well. These would all transform the ball's path in space.

OpenStudy (anonymous):

Thank you

OpenStudy (accessdenied):

You're welcome!